OpenStudy (anonymous):

Two trains leave stations 392 miles apart at the same time and travel toward each other. One train travels at 65 miles per hour while the other travels at 75 miles per hour. How long will it take for the two trains to meet?

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (anonymous):

Closing speed is (75+65) = 140 mph Closing time = distance/rate = 392/140 = 2.8 hrs = 2hr 48minutes

OpenStudy (anonymous):

So is the answer 2 hours and 48 minutes?!